# Safety-Driven Response Adaptive Randomisation: An Application in Non-inferiority Oncology Trials

### v1

@@ 2.3 Allocation Proportion Target
In our simulation design, we allow for $u=1,…,U$ update times for $\hat{\pi}_{E}$, which occur independently of the $k$-analyses defined by the group-sequential design (where interim efficacy analyses are assumed to occur only once, at $k=1$).
Specifically, the randomisation proportion $\hat{\pi}_{E}$ is updated sequentially:
  • For patients enrolled before the first update time ($u=1$), $\hat{\pi}_{E}$ is set to 0.5. This is the burn-in stage of the RAR feature of this design [26].
@@
The number of update times ($U$) and the corresponding calendar occasions should be specified in the study protocol. The updates may be equally spaced across the enrolment period or not, depending on the design objectives. However, in the present design, we did not investigate how varying the number and timing of updates affects the design characteristics. Instead, we fixed equally spaced updates every three months, starting from month 3 through month 45, assuming a total enrolment period of 48 months. Crucially, each $\hat{\pi}_{E}$ is estimated using all available data up to a specific update time.
2.3.1 Elastic Adaptation of Allocation Proportion: Integration of Efficacy and Safety
A final element involves a function designed to modulate the RAR adaptation based on safety, utilizing the estimated association between safety and efficacy endpoints. The underlying idea is that if there’s sufficient evidence of non-inferiority, we would favour allocating more patients to the better-tolerated arms. Conversely, if non-inferiority is not supported by the current evidence in the study, the adaptation would shift away from those arms.
This is achieved by the inclusion in the design of an elastic function that dynamically adjusts the estimated allocation target proportion $\hat{\pi}_{E}$ based on the observed treatment effect on PFS, during the enrolment period. The SAFER function is defined in equation (9).
@@ 2.3.1 Elastic Adaptation of Allocation Proportion: Integration of Efficacy and Safety
$SAFER(\hat{\pi}_{E})=\begin{cases}0.5&\text{if }\hat{\Phi}\leq 0.5\\ 0.5+(\hat{\pi}_{E}-0.5)\cdot\left(1-\left(1-\frac{\hat{\Phi}-0.5}{0.5}\right)^{\eta}\right)&\text{if }0.5<\hat{\Phi}<1\\ \hat{\pi}_{E}&\text{if }\hat{\Phi}=1\end{cases}$ (9)
@@
Note that the maximum possible adjusted allocation target to be used in the design is equal to the standard adaptive allocation proportion $\hat{\pi}_{E}$ and the minimum adjusted allocation proportion is set to 0.5, representing equal allocation. The intermediate cases are introduced through a tuning parameter $\hat{\Phi}$ that quantifies the association between the safety and efficacy endpoints. Specifically, $\hat{\Phi}$ is estimated using the cumulative density function (CDF) of the absolute z-value, derived from a survival regression model, which corresponds to the estimated treatment effect on PFS. We only consider $\hat{\Phi}\geq 0.5$, effectively excluding the possibility of a negative association between treatment and PFS. Moreover, the parameter $\eta\geq 1$ is the elastic parameter which controls the curvature of the function: $\eta=1$ corresponds to a linear function, while $\eta>1$ results in a concave shape, as illustrated in Figure 1. Although convex shapes ($\eta<1$) are theoretically allowed, we did not explore the values of $\eta<1$, as the linear function is already considered a conservative choice to adjust the allocation proportion in cases of small $\hat{\Phi}$ values.
@@
Figure 1: Target function of the SAFER design for two target allocations, $\hat{\pi}_{E}=0.7$ and $\hat{\pi}_{E}=0.9$, each evaluated under three shape parameters $\eta=1,2,5$. The SAFER function increases monotonically in $\Phi_{z}$ and reflects more aggressive allocation as both $\hat{\pi}_{E}$ and $\eta$ increase.
3 Performance Evaluation Through Simulation: redesigning a non-inferiority study
@@
This section details the extensive simulation studies conducted to systematically assess the performance of the proposed SAFER design methodology. These simulations aim to evaluate key operating characteristics including statistical power and allocation patterns under various scenarios. The findings presented herein are based on a specific case study, which we redesign using SAFER and help us provide a comprehensive understanding of the SAFER’s design features and its potential advantages in different trial settings.
@@ 3.1 Motivating example: a follow-up to the CAPP-IT study
In the following subsections, we present a re-design of this study using our SAFER approach. Specifically, we formulate a non-inferiority hypothesis on the primary efficacy endpoint (PFS), while for the safety-driven RAR, we focus on the time to dose reduction or drug discontinuation due to capecitabine-induced side effects as the safety endpoint. Moreover, we assume a fixed total duration of the trial equal to 60 months, considering a uniformly distributed enrolment of 48 months and 12 months of follow-up. Assuming a monthly event rate ($\lambda_{C}$) of 0.069, which corresponds to a median PFS of approximately 10 months, and a $HR_{H0}$ of 1.25 ($\lambda_{E}/\lambda_{C}$), the median PFS in the experimental arm must exceed 8 months to demonstrate non-inferiority with respect to the control arm.
3.2 Modelling efficacy and safety endpoints association
To thoroughly assess SAFER’s performance, understanding the interplay between primary efficacy and safety/tolerability endpoints is essential. This section details how this relationship is incorporated into our framework, specifically addressing scenarios where these endpoints are not assumed to be independent.
@@ 3.3 Metrics to assess performance
Through simulations, we evaluated six distinct scenarios to quantify the trial’s performance under various conditions (described below), providing insights into its statistical properties and broader implications. For each scenario, we conducted 10,000 iterations, ensuring a Monte Carlo error of less than 0.5% for both power and type I error rate [28].
We report the overall study power, calculated as the proportion of simulated trials that rejected $H_{0}$ under the alternative hypothesis ($H_{1}$), aggregated across all simulation studies. Moreover, the estimated allocation proportion to the experimental arm is defined as the average proportion of patients assigned to treatment $E$ at the end of each simulated trial.
To assess the performance of the design and compare the SAFER approach with the fixed allocation proportion approach, we provide benchmark results—specifically, power and allocation proportions—obtained from a fixed randomisation design under various allocation proportions (from 0.5 to 0.8). These benchmarks were derived using a theoretical (non-simulation-based) approach, based on the normal approximation of the outcome as defined in equations (5) and (6), and are presented in Figure 3. The power loss is attributed to changes in the allocation proportion from 0.5 to 0.8. The overall power is maximized by $\pi_{E}=0.5$ since the $\lambda_{C}$ and $\lambda_{E}$ are assumed equal, and decreases to approximately 0.79 when $\pi_{E}=0.6$, 0.74 when $\pi_{E}=0.7$, and 0.63 when $\pi_{E}=0.8$.
Figure 3: Theoretical loss of power, using normal approximation, for the primary endpoint analysis corresponding to diverse Information Fractions of 50% and allocation proportions $\pi_{E}$, from 0.5 to 0.8.
@@
Furthermore, to empirically verify the validity of the standard regularity conditions, we conducted a simulation using exponential distributions for survival outcomes, varying the information fraction (IF) at interim from 0.2 to 0.5; the results are reported in Table 1. The observed loss of power at increasing allocation proportions closely aligns with the theoretical results in Figure 3, suggesting that for further comparisons, the normal approximation can be reliably used in this setting. Interestingly, it seems that the IF does not modify the observed loss of power.
Table 1: Power using exponential distributed outcomes with fixed allocation proportion and varying information fraction at interim analysis.

### v2

@@ 2.3 Allocation proportion
It is worth highlighting that we define $\pi_{E}$ as the true target allocation proportion, computed by plugging the true values of $\theta_{E}^{(s)}$ and $\theta_{C}^{(s)}$ into equation 8, in accordance with the simulation settings used for safety data generation. In contrast, $\hat{\pi}_{E}$ denotes the estimated target allocation proportion, obtained by substituting $\hat{\theta}_{E}^{(s)}$ and $\hat{\theta}_{C}^{(s)}$ i.e., the estimations of the respective parameters, into the same equation.
Specifically, in our simulation design, we allow for $u=1,\ldots,U$ update times for $\hat{\pi}_{E}$, which occur independently of the $k$-analyses defined by the group-sequential design (where interim efficacy analyses are assumed to occur only once, at $k=1$). Therefore, the randomisation proportion $\hat{\pi}_{E}$ is updated sequentially:
  • For patients enrolled before the first update time ($u=1$), $\hat{\pi}_{E}$ is set to 0.5. This is the burn-in stage of the RAR feature of this design [26, 27].
@@
The number of update times ($U$) and the corresponding calendar occasions should be specified in the study protocol. The updates may be equally spaced across the enrolment period or not, depending on the design objectives and practical considerations. We fixed equally spaced updates every three months, starting from month 3 through month 45, assuming a total enrolment period of 48 months. Crucially, each $\hat{\pi}_{E}$ is estimated using all available data up to a specific update time.
2.3.1 Efficacy and safety association
To incorporate safety into RAR, the design includes a modulation function that dynamically combines emerging evidence on safety and efficacy. When the data indicate that the experimental treatment is better tolerated and that this improved safety profile translates into a favourable efficacy outcome, the algorithm increases the probability of assigning future patients to the experimental arm. If the data do not confirm an efficacy benefit despite the safety advantage, randomisation resets to an equal (1:1) allocation between the two arms. This adaptive mechanism is governed by a parameter $\hat{\Phi}$, which quantifies the strength of the evidence supporting the impact of improved safety on efficacy. It is defined as the cumulative distribution function (CDF) of the $z$-value associated with the estimated treatment effect on PFS, obtained from the survival model. In fact, a simple correlation metric between safety and efficacy is not appropriate in this context, as it fails to account for the specific characteristics of the survival endpoint. In particular, both endpoints are subject to truncation: the time-to-event safety endpoint (e.g., time to dose reduction or drug discontinuation) and the PFS endpoint are both affected by administrative censoring. Moreover, the safety endpoint is inherently truncated by the occurrence of PFS events, since once progression occurs, the follow-up ends. As a result, the two endpoints are structurally correlated, and conventional correlation coefficients may yield misleading estimates of their association. Specifically, let:
  • $\hat{\theta}_{\text{PFS}}$ denote the estimated treatment effect (e.g., log hazard ratio) on PFS;
  • $\text{SE}(\hat{\theta}_{\text{PFS}})$ its standard error;
  • $Z=\dfrac{\hat{\theta}_{\text{PFS}}}{\text{SE}(\hat{\theta}_{\text{PFS}})}$ the corresponding $z$-statistic;
  • $\Phi(\cdot)$ the standard normal cumulative distribution function.
Since we expect a hazard ratio lower than 1 in case of treatment effect, we take the absolute value of the z-statistic. Therefore, the adaptation weight is defined as:
$\hat{\Phi}=\Phi\left(|Z|\right)=\Phi\left(\left|\dfrac{\hat{\theta}_{\text{PFS}}}{\text{SE}(\hat{\theta}_{\text{PFS}})}\right|\right)$ (9)
Specifically, this parameter $\hat{\Phi}$ is incorporated into the allocation function as follows:
  • When the experimental arm is better tolerated (i.e., the estimated Neyman allocation is greater than 0.5) and $\hat{\Phi}$ exceeds a predefined threshold, the allocation probability for the experimental arm is increased.
  • When $\hat{\Phi}$ is below the threshold, suggesting insufficient evidence, allocation probabilities revert toward equal randomisation.
Since the $\hat{\Phi}$ ranges from 0 to 1, the default threshold is set at 0.5 (corresponding to a z-statistic equal to zero). However, sensitivity analyses should be conducted to assess the effect of alternative choices. Therefore, the SAFER function is defined as:
@@ 2.3.1 Efficacy and safety association
$\text{SAFER}(\hat{\pi}_{E})=\begin{cases}0.5&\text{if }\hat{\Phi}\leq 0.5\\ 0.5+(\hat{\pi}_{E}-0.5)\cdot\left(1-\left(1-\dfrac{\hat{\Phi}-0.5}{0.5}\right)^{\eta}\right)&\text{if }0.5<\hat{\Phi}<1\\ \hat{\pi}_{E}&\text{if }\hat{\Phi}=1\end{cases}$ (10)
@@
Note that the maximum allocation proportion used in the SAFER design corresponds to the estimated Neyman allocation $\hat{\pi}_{E}$ when $\hat{\Phi}=1$, while the minimum is fixed at 0.5, representing complete randomisation, when $\hat{\Phi}\leq 0.5$. Sensitivity analyses should also explore scenarios with upper bounds below 1 (e.g., 0.9) to assess robustness. Moreover, the parameter $\eta$ is the elastic parameter which controls the curvature of the function: $\eta=1$ corresponds to a linear function, while $\eta>1$ results in a concave shape, as illustrated in Figure 1. Although convex shapes ($\eta<1$) are theoretically allowed, we did not explore the values of $\eta<1$, as the linear function is already considered a conservative choice to adjust the allocation proportion in cases of small $\hat{\Phi}$ values.
@@
Figure 1: Target function of the SAFER design for two target allocations, $\hat{\pi}_{E}=0.7$ and $\hat{\pi}_{E}=0.9$, each evaluated under three shape parameters $\eta=1,2,5$. The SAFER function increases monotonically in $\hat{\Phi}$ and reflects more aggressive allocation as both $\hat{\pi}_{E}$ and $\eta$ increase.
3 SAFER design evaluation
@@
This section details the extensive simulation studies conducted to systematically assess the performance of the proposed SAFER design methodology. These simulations aim to evaluate key operating characteristics including statistical power and the estimated allocation proportion under various scenarios. The findings presented herein are based on a specific case study, which we redesign using SAFER and help us provide a comprehensive understanding of the SAFER’s design features and its potential advantages in different trial settings.
@@ 3.1 Motivating example: a follow-up to the CAPP-IT study
In the following subsections, we present a re-design of this study using our SAFER approach. Specifically, we formulate a non-inferiority hypothesis on the primary efficacy endpoint (PFS), while for the safety-driven RAR, we focus on the time to dose reduction or drug discontinuation due to capecitabine-induced side effects as the safety endpoint. Moreover, we assume a fixed total duration of the trial equal to 60 months, considering a uniformly distributed enrolment of 48 months and 12 months of follow-up. Assuming a monthly event rate in the control arm ($\lambda_{C}$) of 0.069, which corresponds to a median PFS of approximately 10 months, and a $HR_{H0}$ of 1.25 ($\lambda_{E}/\lambda_{C}$), the median PFS in the experimental arm must exceed 8 months to demonstrate non-inferiority with respect to the control arm.
3.2 Modelling the association between safety and efficacy endpoints
@@ 3.3 Metrics to assess performance
We define a safety-driven RAR design as one in which the allocation proportion is updated solely based on the estimated Neyman allocation proportion. In contrast, we refer to the SAFER design when the allocation updates incorporate the SAFER targeting function, which accounts for the observed association between efficacy and safety.
To compare these adaptive designs with a fixed allocation approach, we provide benchmark results focusing on statistical power. Specifically, we assess power using both a theoretical (non-simulation-based) approach as defined in equations (5) and (6) (see Figure A1) and a simulation-based approach, which employs exponentially distributed survival times as outcomes. The latter allows us to empirically assess the validity of the asymptotic assumptions underlying the theoretical approximation (see Table A1).
@@
Specifically, in Figure A1 the overall power is maximized by $\pi_{E}=0.5$ and decreases to approximately 0.79 when $\pi_{E}=0.6$, 0.74 when $\pi_{E}=0.7$, and 0.63 when $\pi_{E}=0.8$. In Table A1, the observed loss of power at increasing allocation proportions closely aligns with the theoretical results in Figure A1, suggesting that for further comparisons, the normal approximation can be reliably used in this setting. Interestingly, it seems that the Information Fraction (IF) at interim does not modify the observed loss of power.
Moreover, through simulations, we evaluate several distinct scenarios to assess the design under various conditions (described below), providing insights into its statistical properties and broader implications. As metrics, we report the overall study power and type-I-error rate, calculated as the proportion of simulated trials that rejected $H_{0}$ under $H_{1}$ and $H_{0}$, respectively, the $power_{1}$, calculated as the proportion of simulated trials that rejected $H_{0}$ under $H_{1}$ at interim analysis (probability of early efficacy stop), the estimated allocation proportion to the experimental arm, defined as the average proportion of patients assigned to treatment $E$ at the end of each simulated trial, and the AE rate per arm calculated as the total number of AEs divided by the total person-years, i.e., the number of patients multiplied by the follow-up duration in years, to yield a rate per patient-year. For each scenario, we conducted 10,000 iterations, ensuring a Monte Carlo error of less than 0.5% for both power and type-I error rate control [29].
3.4 Simulating scenarios
In this section, we will describe in detail the distinct scenarios used for the simulation study.
Scenario 0: Safety-driven RAR design (without a SAFER approach) under independence between the efficacy and safety endpoints.
This scenario assesses the impact of a safety-driven RAR design—without implementing the SAFER approach—on power and type-I error rate control, under the assumption of independence between the efficacy and safety endpoints. The median time to dose reduction or drug discontinuation is simulated at 1.5 months in the control group, while in the experimental group, it is varied across values of 1.5, 2.25, 3.5, and 6 months corresponding to $\pi_{E}$ equal to 0.5, 0.60, 0.70, and 0.80, respectively. Moreover, we set $U=15$, resulting in updating $\hat{\pi}_{E}$ every three months until month 45, following an initial burn-in period during which a 1:1 randomization is applied for the first three months (about 60 patients).
Scenario 1: Complete randomization design under diverse level of association between the efficacy and safety endpoints.
This scenario is designed to assess the impact of diverse level of association between the efficacy and safety endpoints on statistical power and AEs rate. Specifically, we simulate increasing levels of association between the endpoints, with $E[\gamma_{1}]$ set to 0.001, 0.005, 0.01, 0.03, and 0.05, corresponding to very weak, weak, moderate, strong, and very strong associations, respectively. Moreover, we set the median AEs time in each arm corresponding to $\pi_{E}$ from 0.5 (no difference in safety between arms) to 0.8. To evaluate the effect of association between endpoint at the net of RAR, we simulate under a 1:1 allocation ratio (fair coin toss).
Scenario 2: Safety-driven RAR design (without a SAFER approach) under diverse level of association between the efficacy and safety endpoints.
This scenario is designed as a counterpart to scenario 1 and aims to assess the impact on power, allocation proportion, and AEs rate of the safety-driven RAR design. Therefore, at each time point $u=1,\ldots,U$, we compute $\hat{\pi}_{E}=\hat{\theta}_{E}^{(s)}/(\hat{\theta}_{E}^{(s)}+\hat{\theta}_{C}^{(s)})$ and use it as the allocation probability to arm $E$ for the randomization of patients in the subsequent three months. Under this scenario, we expect power to decrease in the presence of weak association between endpoints, and to increase as association strengthens, potentially due to the enhanced effectiveness of the experimental treatment on the efficacy endpoint driven by prolonged well-tolerated drug exposure.
Scenario 3: SAFER design: impact on power under diverse level of association between the efficacy and safety endpoints, varying the $\eta$ values.
This scenario is designed to evaluate the impact on power, allocation proportion, and AEs rate of the SAFER design, by varying the elastic parameter $\eta$ values. Specifically, scenarios 3a and 3b report the results of simulations conducted with $\eta=1$ and $\eta=5$, respectively. These values correspond to the absence of elasticity and a high degree of elasticity in the target SAFER function, as reported in Figure 1.
Scenario 4: Impact of differential timing in the observation of endpoint data on $SAFER(\hat{\pi}_{E})$.
This scenario aims to assess the speed at which $SAFER(\hat{\pi}_{E})$ reaches the target level, when the median times of the safety and efficacy endpoints are either close or far apart i.e., delayed efficacy outcome [10]. For instance, we expect a faster pattern when the median PFS is close to the median time to dose reduction or drug discontinuation because there will be enough PFS events for the $\hat{\Phi}$ parameter estimation, early in the recruitment period. On the other hand, when evidence of the positive association between safety and efficacy emerges later in the recruitment, we expect the SAFER function to down-weight the allocation proportion for a longer period until sufficient evidence on PFS becomes available. Under this scenario, the median PFS ranges from 3 to 24 months.
Scenario 5: Impact of informative drop-out (related to the safety endpoint) on $SAFER(\hat{\pi}_{E})$ and power.
This scenario aims to assess the impact of informative drop-out on the operating characteristics of the SAFER design. We define informative drop-out as a situation in which a patient is observed until experiencing a side effect and subsequently drops out of the study. In practice, this means that the patient is followed until dose reduction or drug discontinuation occurs, after which neither the time to PFS nor the occurrence of the progression event is observed. We simulated the average allocation proportion and power under varying informative drop-out rates, ranging from 5% (very low) to 25% (very high), with drop-outs uniformly distributed over the accrual period. To address the potential bias introduced by informative drop-out, we considered the strategies for handling intercurrent events as defined in the Estimands framework [30]. Specifically, we considered the composite strategy, where the time-to-PFS is defined as the last available follow-up time, with the occurrence of the progression = yes.
Scenario 6: Impact of under-reported safety events on $SAFER(\hat{\pi}_{E})$ and power.
This scenario aims to assess the impact of under-reported safety events on the operating characteristics of the SAFER design. Unlike the informative drop-out phenomenon—where patients are observed until an adverse event occurs and then drop out—in the case of under-reported safety events, it is not recorded at all and only the survival outcome (i.e., progression and the corresponding time-to-PFS) is available [31].
3.5 Simulation Results
Table 1 presents the results of the simulations under Scenario 0 (independence between endpoints). Both the estimated allocation proportion of patients assigned to the experimental arm $N_{E}/N$ and the power closely align with the target allocation $\pi_{E}$ and the theoretical power reported in Table A1, respectively. Minor discrepancies can be attributed to the burn-in period, as Table A1 reports results based on unequal allocation without any burn-in, whereas Scenario 0 includes this initial adjustment phase. The type-I error rate remains close to the nominal value of 0.05, with minor increases. However, these increases do not show a consistent pattern and are likely due to simulation variability. Therefore, neither the safety-driven RAR design nor the information fraction (IF) appears to negatively affect type-I error control under the null hypothesis.
Table 1: Estimated allocation proportion, power, and type-I error rate summary: Scenario 0
Allocation Power/Type-I error rate Information Fraction
$\pi_{E}$ $N_{E}/N$ 0.2 0.3 0.4 0.5
0.5 0.5 Power 0.80 0.80 0.80 0.80
0.5 Type-I error rate 0.048 0.050 0.048 0.053
0.6 0.59 Power 0.78 0.79 0.78 0.78
0.59 Type-I error rate 0.047 0.050 0.055 0.053
0.7 0.69 Power 0.74 0.74 0.73 0.73
0.69 Type-I error rate 0.049 0.051 0.051 0.053
0.8 0.78 Power 0.66 0.66 0.65 0.66
0.78 Type-I error rate 0.050 0.055 0.056 0.051
$\pi_{E}$ = True target allocation value by simulation setting; $N_{E}/N$ = Proportion of patients assigned to the experimental arm using a safety-driven RAR design.
Table 2 presents the results of the simulations under Scenarios 1, 2, 3a, and 3b. Overall, it shows how the SAFER method effectively balances the benefits of a safety-driven RAR design with the need to maintain statistical validity. Scenario 1 is simulated using a complete randomization approach. Therefore, the observed increase in both $power_{1}$ and overall power can be attributed to the influence of the safety endpoint on the efficacy endpoint. Specifically, the longer the time to AEs in arm E, the longer the PFS tends to be in the same arm, which increases the probability of rejecting the null hypothesis. This trend is amplified by the strength of the association between the endpoints, with $power_{1}$ reaching 0.82 when $\pi_{E}=0.8$ and the association is very strong.
Scenario 2 is simulated using a safety-driven RAR design, without the SAFER approach. As the estimated allocation proportion $N_{E}/N$ increases, both $power_{1}$ and the overall power decrease in the presence of very weak or weak associations between endpoints. Conversely, when the strength of association increases, a gain in power is observed. This improvement is driven by the meaningful positive effect on PFS resulting from drug exposure free from AEs, as described in Section 3.2.
Under Scenario 3a, the SAFER design is implemented with the elastic parameter $\eta$ fixed at 1. In the presence of weak endpoint associations, the SAFER approach effectively shrinks the estimated allocation proportion towards 0.5, thereby restoring statistical power while still favouring the better-tolerated arm—though to a lesser extent. Conversely, when the association between endpoints is strong, the design allows higher allocation proportions in favour of the experimental arm without compromising statistical power. A similar pattern is observed in Scenario 3b, where the higher value of the elastic parameter $\eta=5$ allows for a more aggressive approach in reaching the target allocation proportion. This comes at the cost of a slight reduction in power compared to Scenario 3a.
A crucial metric for evaluating the benefit of the proposed approach is the reduction in the AE rate per patient-year. Specifically, under the CR design (Scenario 1), the AE rate decreases as $\pi_{E}$ increases, from 1.28 when $\pi_{E}=0.5$ to approximately 1.05 when $\pi_{E}=0.8$. However, thanks to the safety-driven RAR design in Scenario 2 and the SAFER design in Scenarios 3a and 3b, the probability of being assigned to the better-tolerated arm increases. As a result, a larger reduction in the AE rate per patient-year is observed, with values dropping below 1 for high $N_{E}/N$ values and strong endpoint associations.
Table 2: Estimated allocation proportion, power, and AEs rate summary: Scenarios 1, 2, 3a, and 3b (IF = 0.5)